Given the following demand and supply functions: Supply: Os-12+1p Demand: Qd 84-3p 0.02Y, where Y Consumer income per month Solve for the following given consumer income is $5,000/mo Equilibrium Price : $Li (round your calculation to the nearest penny).

Round to two decimal points.
Answer #1

Solution -

QS= -12 + P

Qd=84 - 3P + 0.02Y

The Y=$5000 per month is given

Qd=84-3P+0.02Y

=84-3P+0.02(5000)

=84-3P+100

=184-3P

The equilibrium condition

Qd=Qs

184-3P= -12+P

3P+P =184+12

4P =196

P =196/4

P =$49

Hence the equilibrium price is  $49.